OpenStudy (anonymous):

2[3+2(x-6)]+3{-2(x-5)+9]

OpenStudy (anonymous):

PEMDAS ? :)

OpenStudy (anonymous):

-2x+39 not sure though

OpenStudy (anonymous):

-2x + 27 mine :)

OpenStudy (anonymous):

its 2x+39, i have the answer but not sure how you got it??

OpenStudy (anonymous):

yeah right its 2x + 39 my mistake :3

OpenStudy (triciaal):

do the inside parenthesis before the outside I think you missed a close parenthesis after the 5 simplify the inside results as you go along @Shikatanori do you want to show your work since you already did it?

OpenStudy (anonymous):

1.) 2[3+2(x-6)] + 2.) 3{-2(x-5)+9] ? = as i said before PEMDAS ? :) first parentesis ? . 1.) fist equation 2(x-6) = 2x-12+3=2[2x-9] = 4x-18 then 2.) second equation -2(x-5)= -2x+10+9 = 3[-2x+19] = -6x+57 then thats it 4x-18-6x+57 = -2x+39 ? :3 Sorry my mistake HEhe

OpenStudy (anonymous):

how did you get the answer

OpenStudy (anonymous):